# Effective Medium Theory for Time-modulated Subwavelength Resonators

### v1

@@ 3.1 Transfer and Scattering Matrices
Consider the resonator $D_{i}$, and let the wave field on the left- and right-hand sides of $D_{i}$ be given by $v(x)=\alpha^{i}\mathrm{e}^{\mathrm{i}kx}+\beta^{i}\mathrm{e}^{-\mathrm{i}kx}$ and $v(x)=\alpha^{i+1}\mathrm{e}^{\mathrm{i}kx}+\beta^{i+1}\mathrm{e}^{-\mathrm{i}kx}$, respectively. Then the transfer matrix $\tilde{S}_{i}$ corresponding to $D_{i}$ is given by
@@
If we let the interior wave field be given by $v(x)=a\mathrm{e}^{\mathrm{i}k_{\mathrm{r}}x}+b\mathrm{e}^{-\mathrm{i}k_{\mathrm{r}}x},\ x\in D_{i}$, the continuity and transmission conditions give us
@@
By plugging in (3.8) into (3.3), we obtain
$\begin{bmatrix}\tilde{\alpha}^{i+1}\\ \tilde{\beta}^{i}\end{bmatrix}=g\begin{bmatrix}1&\mathrm{e}^{-2\mathrm{i}kz_{i}}\\ \mathrm{e}^{2\mathrm{i}kz_{i}}&1\end{bmatrix}\begin{bmatrix}\alpha^{i}\\ \beta^{i+1}\end{bmatrix},$ (3.9)
@@
where $v^{\mathrm{in}}(x):=\alpha^{1}\mathrm{e}^{\mathrm{i}kx}+\beta^{N+1}\mathrm{e}^{-\mathrm{i}kx}$ is the overall incident wave on the system of $N$ resonators. Moreover, we denote the incident wave field on $D_{i}$ by $v^{\mathrm{in}}_{i}(x)$ and the scattered wave field of $D_{i}$ by $v^{\mathrm{sc}}_{i}(x)$. The incident and scattered wave field of $D_{i}$ are given by
@@ 3.2 Homogenised Equation
Equation (3.15) is a linear system of equations: define the vectors $\mathbf{x},\,\mathbf{b}\in\mathbb{C}^{N}$ and the matrix $T\in\mathbb{C}^{N\times N}$ with entries
@@
$\mathbf{x}-T\mathbf{x}=\mathbf{b}.$ (3.17)
@@
Let $A\subset\mathbb{R}$ be a measurable set, we define
@@
We define the $L^{\infty}$-function $V(x):=\beta\Lambda\tilde{V}(x)$. The following assumption is feasible based on [5, Lemma 4.1].
@@
We assume that for any $f\in C^{0,\alpha}(\mathcal{U})$
@@
for some constants $10<\alpha<$ 1 and $C>0$.
@@
Analogously as in [5], this operator can be seen as the continuum limit of the matrix operator $T$. Since we define $\mathcal{T}$ in an equivalent manner as in the higher-dimensional case, the properties proven in [5, Lemma 4.2] still hold true in one dimension.
@@
which is exactly the homogenised equation. Similarly as in [5], it can be proven that $\psi$ is the continuum limit of the solution $\mathbf{x}$ to (3.17).
@@ 4 Time-dependent Metamaterial
$\begin{cases}\alpha^{i}_{n}\mathrm{e}^{\mathrm{i}k^{(n)}x_{i}^{-}}+\beta^{i}_{n}\mathrm{e}^{-\mathrm{i}k^{(n)}x_{i}^{-}}=\sum\limits_{j=-K}^{K}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x_{i}^{-}}+b_{j}^{1}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x_{i}^{-}}\right)f_{n}^{j,i},\\ \alpha^{i+1}_{n}\mathrm{e}^{\mathrm{i}k^{(n)}x_{i}^{+}}+\beta^{i+1}_{n}\mathrm{e}^{-\mathrm{i}k^{(n)}x_{i}^{+}}=\sum\limits_{j=-K}^{K}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x_{i}^{+}}+b_{j}^{1}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x_{i}^{+}}\right)f_{n}^{j,i},\end{cases}$ (4.1)
$\begin{cases}\delta k^{(n)}\left(\alpha^{i}_{n}\mathrm{e}^{\mathrm{i}k^{(n)}x_{i}^{-}}-\beta^{i}_{n}\mathrm{e}^{-\mathrm{i}k^{(n)}x_{i}^{-}}\right)=\sum\limits_{j=-K}^{K}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x_{i}^{-}}-b_{j}^{1}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x_{i}^{-}}\right)\lambda_{j}^{i}f_{n}^{j,i},\\ \delta k^{(n)}\left(\alpha^{i+1}_{n}\mathrm{e}^{\mathrm{i}k^{(n)}x_{i}^{+}}-\beta^{i+1}_{n}\mathrm{e}^{-\mathrm{i}k^{(n)}x_{i}^{+}}\right)=\sum\limits_{j=-K}^{K}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x_{i}^{+}}-b_{j}^{1}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x_{i}^{+}}\right)\lambda_{j}^{i}f_{n}^{j,i},\end{cases}$ (4.2)
@@
$\boldsymbol{\alpha}^{i}:=\begin{bmatrix}\alpha^{i}_{-K}\\ \vdots\\ \alpha_{K}^{i}\end{bmatrix},\quad\boldsymbol{\beta}^{i}:=\begin{bmatrix}\beta^{i}_{-K}\\ \vdots\\ \beta^{i}_{K}\end{bmatrix},\quad\mathbf{w}^{i}:=\begin{bmatrix}\boldsymbol{\alpha}^{i}\\ \boldsymbol{\beta}^{i}\end{bmatrix},\quad\mathbf{v}_{j}^{i}:=\begin{bmatrix}a_{j}^{i}\\ b_{j}^{i}\end{bmatrix},\quad\mathbf{v}^{i}:=\begin{bmatrix}\mathbf{v}_{-K}^{i}\\ \vdots\\ \mathbf{v}_{K}^{i}\end{bmatrix}.$ (4.3)
@@
$E_{\pm}:=\mathrm{diag}\left(\mathrm{e}^{\pm\mathrm{i}k^{(n)}x^{+}}\right)_{n=-K}^{K},\quad K:=\mathrm{diag}\left(k^{(n)}\right)_{n=-K}^{K},$ (4.4)
@@
$\mathcal{M}^{-}:=\begin{bmatrix}GE_{-}&GE_{+}\\ \delta GKE_{-}&-\delta GKE_{+}\end{bmatrix},\quad\mathcal{T}^{-}:=\begin{bmatrix}T^{-}\\ T^{-}_{\lambda}\end{bmatrix},\quad\mathcal{M}^{+}:=\begin{bmatrix}E_{+}&E_{-}\\ \delta KE_{+}&-\delta KE_{-}\end{bmatrix},\quad\mathcal{T}^{+}:=\begin{bmatrix}FT^{+}\\ FT^{+}_{\lambda}\end{bmatrix}.$ (4.6)
@@
$\mathcal{M}^{-}\mathbf{w}^{i}=\mathcal{T}\mathbf{v}^{i},\quad\mathcal{M}^{+}\mathbf{w}^{i+1}=\mathcal{T}^{+}\mathbf{v}^{i}.$ (4.7)
We have now transformed the four equations (4.1) and (4.2) that are true for each $n$ into two linear systems of equations, which contain all the modes $n$. By combining the two equations into a single one, we arrive at the following proposition, which defines the time-dependent scattering matrix. We assume that $\omega$ is not a multiple of $\Omega$, whereby a direct calculation shows that $\mathcal{M}^{+}$ and $\mathcal{T}^{-}$ are invertible.
@@
$\mathbf{w}^{i+1}=\tilde{S}_{i}\mathbf{w}^{i},\quad\tilde{S}_{i}:=\left(\mathcal{M}^{+}\right)^{-1}\mathcal{T}^{+}\left(\mathcal{T}^{-}\right)^{-1}\mathcal{M}^{-}=:\begin{bmatrix}\tilde{S}_{11}&\tilde{S}_{12}\\ \tilde{S}_{21}&\tilde{S}_{22}\end{bmatrix}\in\mathbb{C}^{2(2K+1)\times 2(2K+1)},$ (4.8)
@@
As in the previous section, we proceed by taking the limit $\ell\to 0$ and derive a point-scattering approximation equivalent to Lemma 3.2. We take the same asymptotic scaling as before, whereby $\delta=\gamma\ell,\,\omega=\mu\ell,\,\Omega=\xi\ell,$ and $\lambda_{j}^{i}=c_{j}^{i}\ell$, for fixed $\gamma,\,\mu,\,\xi>0$ independent of $\ell$. Here, recall that $\lambda_{j}^{i}$ are the wave numbers in the $i$th resonator, as in 2.1. Note that the exact definition of $c_{j}^{i}$ depends on $\mu$ and $\Omega$. Although there is no closed-form expression for these coefficients, they can be efficiently computed as the eigenvalues of the temporal Sturm-Liouville problem in the interior of $D_{i}$ (see, for example, [1, Lemma III.3] or [13, Section V]).
@@
$S_{\mathcal{M}^{+}}=-\delta KE_{-}-(\delta KE_{+})\left(E_{+}\right)^{-1}E_{-}=-2\delta KE_{-}\quad\implies\quad S_{\mathcal{M}^{+}}^{-1}=-\frac{1}{2\delta}E_{+}K^{-1}.$ (4.15)
@@
$\left(\mathcal{M}^{+}\right)^{-1}=\frac{1}{2}\begin{bmatrix}E_{-}&\frac{1}{\delta}E_{-}K^{-1}\\ E_{+}&-\frac{1}{\delta}E_{+}K^{-1}\end{bmatrix}.$ (4.16)
@@
$\mathcal{M}^{-}=\underbrace{\begin{bmatrix}G&0\\ 0&G\end{bmatrix}}_{=:\mathcal{G}}\begin{bmatrix}E_{-}&E_{+}\\ \delta KE_{-}&-\delta KE_{+}\end{bmatrix}.$ (4.20)
@@
$\tilde{S}_{i}=\frac{1}{2}\begin{bmatrix}E_{-}&\frac{1}{\delta}E_{-}K^{-1}\\ E_{+}&-\frac{1}{\delta}E_{+}K^{-1}\end{bmatrix}\mathcal{F}\begin{bmatrix}T^{+}\hat{T}&T^{+}\hat{T}_{\lambda}\\ T^{+}_{\lambda}\hat{T}&T^{+}_{\lambda}\hat{T}_{\lambda}\end{bmatrix}\mathcal{G}\begin{bmatrix}E_{-}&E_{+}\\ \delta KE_{-}&-\delta KE_{+}\end{bmatrix},$ (4.21)
@@
$\tilde{S}_{11}=\frac{1}{2}\left(\left(E_{-}F(T^{+}\hat{T})+\frac{1}{\delta}E_{-}K^{-1}F(T^{+}_{\lambda}\hat{T})\right)GE_{-}+\left(\delta E_{-}F(T^{+}\hat{T}_{\lambda})+E_{-}K^{-1}F(T^{+}_{\lambda}\hat{T}_{\lambda})\right)GKE_{-}\right),$
$\tilde{S}_{12}=\frac{1}{2}\left(\left(E_{-}F(T^{+}\hat{T})+\frac{1}{\delta}E_{-}K^{-1}F(T^{+}_{\lambda}\hat{T})\right)GE_{+}-\left(\delta E_{-}F(T^{+}\hat{T}_{\lambda})+E_{-}K^{-1}F(T^{+}_{\lambda}\hat{T}_{\lambda})\right)GKE_{+}\right),$
$\tilde{S}_{21}=\frac{1}{2}\left(\left(E_{+}F(T^{+}\hat{T})-\frac{1}{\delta}E_{+}K^{-1}F(T^{+}_{\lambda}\hat{T})\right)GE_{-}+\left(\delta E_{+}F(T^{+}\hat{T}_{\lambda})-E_{+}K^{-1}F(T^{+}_{\lambda}\hat{T}_{\lambda})\right)GKE_{-}\right),$
$\tilde{S}_{22}=\frac{1}{2}\left(\left(E_{+}F(T^{+}\hat{T})-\frac{1}{\delta}E_{+}K^{-1}F(T^{+}_{\lambda}\hat{T})\right)GE_{+}-\left(\delta E_{+}F(T^{+}\hat{T}_{\lambda})-E_{+}K^{-1}F(T^{+}_{\lambda}\hat{T}_{\lambda})\right)GKE_{+}\right).$
@@
where $\tilde{g}^{i}_{nm}:=2\mathrm{i}k^{(n)}g^{i}_{nm}$ and we use the notation $g^{i}:=\left(g^{i}_{nm}\right)_{m,n=-K}^{K}$. Here, $v_{i,m}^{\mathrm{in}}$ is the field which is incident to $D_{i}$, and the scattered field is defined piecewise by

### v3

@@ 3.1 Transfer and Scattering Matrices
Consider the resonator $D_{i}$, and let the wave field on the left- and right-hand sides of $D_{i}$ be given by $v^{\mathrm{sc}}(x)=\alpha^{i}\mathrm{e}^{\mathrm{i}kx}+\beta^{i}\mathrm{e}^{-\mathrm{i}kx}$ and $v^{\mathrm{sc}}(x)=\alpha^{i+1}\mathrm{e}^{\mathrm{i}kx}+\beta^{i+1}\mathrm{e}^{-\mathrm{i}kx}$, respectively. Then the transfer matrix $\tilde{S}_{i}$ corresponding to $D_{i}$ is given by
@@
If we let the interior wave field be given by $v^{\mathrm{sc}}(x)=a\mathrm{e}^{\mathrm{i}k_{\mathrm{r}}x}+b\mathrm{e}^{-\mathrm{i}k_{\mathrm{r}}x},\ x\in D_{i}$, the continuity and transmission conditions give us
@@
By plugging in (3.8) into (3.3), we obtain for sufficiently small $\ell>0$
$\begin{bmatrix}\tilde{\alpha}^{i+1}\\ \tilde{\beta}^{i}\end{bmatrix}=g\begin{bmatrix}1&\mathrm{e}^{-2\mathrm{i}kz_{i}}\\ \mathrm{e}^{2\mathrm{i}kz_{i}}&1\end{bmatrix}\begin{bmatrix}\alpha^{i}\\ \beta^{i+1}\end{bmatrix}+O(\ell^{2}),$ (3.9)
@@
where $v^{\mathrm{in}}(x):=\alpha^{1}\mathrm{e}^{\mathrm{i}kx}+\beta^{N+1}\mathrm{e}^{-\mathrm{i}kx}$ is the overall incident wave on the system of $N$ resonators. Moreover, we denote the incident wave field on $D_{i}$ by $v^{\mathrm{in}}_{i}(x)$ and the scattered wave field of $D_{i}$ by $v^{\mathrm{sc}}_{i}(x)$. Here, $v^{\mathrm{in}}$ is the overall incident field, while $v^{\mathrm{in}}_{i}$ is the field impinging on the $i$-th resonator. In other words, $v^{\mathrm{in}}_{i}$ is given by the sum of the overall incident field and the scattered fields from the other resonators. The incident and scattered wave field of $D_{i}$ are given by
@@ 3.2 Homogenised Equation
Equation (3.15) is a linear system of equations, i.e., we can define the vectors $\boldsymbol{x},\,\boldsymbol{b}\in\mathbb{C}^{N}$ and the matrix $T\in\mathbb{C}^{N\times N}$ with entries
@@
$\boldsymbol{x}-T\boldsymbol{x}=\boldsymbol{b}.$ (3.17)
@@
Let $A\subset\mathcal{U}$ be a measurable set, we define
@@
We define the $L^{\infty}$-function
$V(x):=\left\{\begin{array}[]{l}\beta\Lambda\tilde{V}(x),\quad x\in\mathcal{U},\\ 0,\quad x\in\mathbb{R}\setminus\mathcal{U}.\end{array}\right.$
The following assumption on the regularity of the distribution of the centres of the resonators is feasible based on [7, Lemma 4.1].
@@
Let $k$ be in a fixed neighbourhood of zero. We assume that for any $f\in C^{0,\alpha}(\mathcal{U})$
@@
for some constants $10<\alpha<$ 1 and $C>0$, independent of $k$.
@@
Analogously as in [7], this operator can be seen as the continuum limit of the matrix operator $T$. Since we define $\mathcal{T}$ in an equivalent manner as in the higher-dimensional case, the properties proven in [7, Lemma 4.2] still hold true in one dimension.
@@
which is exactly the homogenised equation. Similarly as in [7], it can be proven that $\psi$ is the continuum limit of the solution $\boldsymbol{x}$ to (3.17).
Remark 3.7.
Note that, in general, (3.24) may not be well-posed since $k^{2}-V$ may change sign inside $\mathcal{U}$. The effective potential $V$ may also be zero in some parts of $\mathcal{U}$ depending on the distribution of the centers of the resonators described by $\tilde{V}$. This makes the study of the existence and uniqueness of a solution to (3.24) a delicate open problem. Moreover, note that if $k^{2}-V$ is negative, then the effective medium is dissipative.
@@ 4 Time-dependent Metamaterial
$\begin{cases}\alpha^{i}_{n}\mathrm{e}^{\mathrm{i}k^{(n)}x_{i}^{-}}+\beta^{i}_{n}\mathrm{e}^{-\mathrm{i}k^{(n)}x_{i}^{-}}=\sum\limits_{j=-K}^{K}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x_{i}^{-}}+b_{j}^{i}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x_{i}^{-}}\right)f_{n}^{j,i},\\ \alpha^{i+1}_{n}\mathrm{e}^{\mathrm{i}k^{(n)}x_{i}^{+}}+\beta^{i+1}_{n}\mathrm{e}^{-\mathrm{i}k^{(n)}x_{i}^{+}}=\sum\limits_{j=-K}^{K}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x_{i}^{+}}+b_{j}^{i}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x_{i}^{+}}\right)f_{n}^{j,i},\end{cases}$ (4.1)
and
$\begin{cases}\delta k^{(n)}\left(\alpha^{i}_{n}\mathrm{e}^{\mathrm{i}k^{(n)}x_{i}^{-}}-\beta^{i}_{n}\mathrm{e}^{-\mathrm{i}k^{(n)}x_{i}^{-}}\right)=\sum\limits_{j=-K}^{K}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x_{i}^{-}}-b_{j}^{i}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x_{i}^{-}}\right)\lambda_{j}^{i}f_{n}^{j,i},\\ \delta k^{(n)}\left(\alpha^{i+1}_{n}\mathrm{e}^{\mathrm{i}k^{(n)}x_{i}^{+}}-\beta^{i+1}_{n}\mathrm{e}^{-\mathrm{i}k^{(n)}x_{i}^{+}}\right)=\sum\limits_{j=-K}^{K}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x_{i}^{+}}-b_{j}^{i}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x_{i}^{+}}\right)\lambda_{j}^{i}f_{n}^{j,i},\end{cases}$ (4.2)
@@
$\boldsymbol{\alpha}^{i}:=\begin{bmatrix}\alpha^{i}_{-K}\\ \vdots\\ \alpha_{K}^{i}\end{bmatrix},\quad\boldsymbol{\beta}^{i}:=\begin{bmatrix}\beta^{i}_{-K}\\ \vdots\\ \beta^{i}_{K}\end{bmatrix},\quad\boldsymbol{w}^{i}:=\begin{bmatrix}\boldsymbol{\alpha}^{i}\\ \boldsymbol{\beta}^{i}\end{bmatrix},\quad\boldsymbol{v}_{j}^{i}:=\begin{bmatrix}a_{j}^{i}\\ b_{j}^{i}\end{bmatrix},\quad\boldsymbol{v}^{i}:=\begin{bmatrix}\boldsymbol{v}_{-K}^{i}\\ \vdots\\ \boldsymbol{v}_{K}^{i}\end{bmatrix}.$ (4.3)
@@
$E_{\pm,\mathrm{l}}:=\mathrm{diag}\left(\mathrm{e}^{\pm\mathrm{i}k^{(n)}x^{-}}\right)_{n=-K}^{K},\,\,E_{\pm,\mathrm{r}}:=\mathrm{diag}\left(\mathrm{e}^{\pm\mathrm{i}k^{(n)}x^{+}}\right)_{n=-K}^{K},\,\,K:=\mathrm{diag}\left(k^{(n)}\right)_{n=-K}^{K},$ (4.4)
@@
$\mathcal{M}^{-}:=\begin{bmatrix}GE_{+,\mathrm{l}}&GE_{-,\mathrm{l}}\\ \delta GKE_{+,\mathrm{l}}&-\delta GKE_{-,\mathrm{l}}\end{bmatrix},\,\,\mathcal{T}^{-}:=\begin{bmatrix}T^{-}\\ T^{-}_{\lambda}\end{bmatrix},\,\,\mathcal{M}^{+}:=\begin{bmatrix}E_{+,\mathrm{r}}&E_{-,\mathrm{r}}\\ \delta KE_{+,\mathrm{r}}&-\delta KE_{-,\mathrm{r}}\end{bmatrix},\,\,\mathcal{T}^{+}:=\begin{bmatrix}FT^{+}\\ FT^{+}_{\lambda}\end{bmatrix}.$ (4.6)
@@
$\mathcal{M}^{-}\boldsymbol{w}^{i}=\mathcal{T}\boldsymbol{v}^{i},\quad\mathcal{M}^{+}\boldsymbol{w}^{i+1}=\mathcal{T}^{+}\boldsymbol{v}^{i}.$ (4.7)
We have now transformed the four equations (4.1) and (4.2), that are true for each $n$, into two linear systems of equations, which contain all the modes $n$. By combining the two equations into a single one, we arrive at the following proposition, which defines the time-dependent scattering matrix. We assume that $\omega$ is not a multiple of $\Omega$, whereby a direct calculation shows that $\mathcal{M}^{+}$ and $\mathcal{T}^{-}$ are invertible.
@@
$\boldsymbol{w}^{i+1}=\tilde{S}_{i}\boldsymbol{w}^{i},\quad\tilde{S}_{i}:=\left(\mathcal{M}^{+}\right)^{-1}\mathcal{T}^{+}\left(\mathcal{T}^{-}\right)^{-1}\mathcal{M}^{-}=:\begin{bmatrix}\tilde{S}_{11}&\tilde{S}_{12}\\ \tilde{S}_{21}&\tilde{S}_{22}\end{bmatrix}\in\mathbb{C}^{2(2K+1)\times 2(2K+1)},$ (4.8)
@@
As in the previous section, we proceed by taking the limit $\ell\to 0$ and derive a point-scattering approximation equivalent to Lemma 3.2. We take the same asymptotic scaling as before:
$\delta=\gamma\ell\quad\omega=\mu\ell,\quad\Omega=\xi\ell$ (4.11)
for fixed $\gamma,\,\mu,\,\xi>0$ independent of $\ell$. Moreover, recall that $\lambda_{j}^{i}$ are the wave numbers in the $i$-th resonator, as in 2.2. The coefficients $\lambda_{j}^{i}$ are given by the eigenvalues of the matrix defined in [2, Lemma III.3] and, under the chosen parameter regime, it is straight-forward to show that $\lambda_{j}^{i}$ is of order $\ell$. As $\ell\to 0$, we then define the coefficients $c_{j}^{i}$ through
$\lambda_{j}^{i}=c_{j}^{i}\ell.$ (4.12)
Note that the exact definition of $c_{j}^{i}$ depends on $\omega$ and $\Omega$. Although there is no closed-form expression for these coefficients, they can be efficiently computed as the eigenvalues of the temporal Sturm-Liouville problem in the interior of $D_{i}$. Following [17, Section V], the coefficients $\lambda_{j}^{i}$ can be phrased as the eigenvalues of a Sturm-Liouville problem in the temporal variable. By taking a Fourier transform of this eigenvalue problem, we obtain the equivalent characterisation given in [2, Lemma III.3]. In the remainder of this paper, we will denote the $n\times n$ identity matrix by $I_{n}$.
@@
$S_{\mathcal{M}^{+}}=-\delta KE_{-,\mathrm{r}}-(\delta KE_{+,\mathrm{r}})\left(E_{+,\mathrm{r}}\right)^{-1}E_{-,\mathrm{r}}=-2\delta KE_{-,\mathrm{r}}\quad\implies\quad S_{\mathcal{M}^{+}}^{-1}=-\frac{1}{2\delta}E_{+,\mathrm{r}}K^{-1}.$ (4.17)
@@
$\left(\mathcal{M}^{+}\right)^{-1}=\frac{1}{2}\begin{bmatrix}E_{-,\mathrm{r}}&\frac{1}{\delta}E_{-,\mathrm{r}}K^{-1}\\ E_{+,\mathrm{r}}&-\frac{1}{\delta}E_{+,\mathrm{r}}K^{-1}\end{bmatrix}.$ (4.18)
@@
$\mathcal{M}^{-}=\underbrace{\begin{bmatrix}G&0\\ 0&G\end{bmatrix}}_{=:\mathcal{G}}\begin{bmatrix}E_{-,\mathrm{l}}&E_{+,\mathrm{l}}\\ \delta KE_{-,\mathrm{l}}&-\delta KE_{+,\mathrm{l}}\end{bmatrix}.$ (4.22)
@@
$\tilde{S}_{i}=\frac{1}{2}\begin{bmatrix}E_{-,\mathrm{r}}&\frac{1}{\delta}E_{-,\mathrm{r}}K^{-1}\\ E_{+,\mathrm{r}}&-\frac{1}{\delta}E_{+,\mathrm{r}}K^{-1}\end{bmatrix}\mathcal{F}\begin{bmatrix}T^{+}\hat{T}&T^{+}\hat{T}_{\lambda}\\ T^{+}_{\lambda}\hat{T}&T^{+}_{\lambda}\hat{T}_{\lambda}\end{bmatrix}\mathcal{G}\begin{bmatrix}E_{-,\mathrm{l}}&E_{+,\mathrm{l}}\\ \delta KE_{-,\mathrm{l}}&-\delta KE_{+,\mathrm{l}}\end{bmatrix},$ (4.23)
@@
$\tilde{S}_{11}=\frac{1}{2}\left(\left(E_{-,\mathrm{r}}F(T^{+}\hat{T})+\frac{1}{\delta}E_{-,\mathrm{r}}K^{-1}F(T^{+}_{\lambda}\hat{T})\right)GE_{+,\mathrm{l}}+\left(\delta E_{-,\mathrm{r}}F(T^{+}\hat{T}_{\lambda})+E_{-,\mathrm{r}}K^{-1}F(T^{+}_{\lambda}\hat{T}_{\lambda})\right)GKE_{+,\mathrm{l}}\right),$
$\tilde{S}_{12}=\frac{1}{2}\left(\left(E_{-,\mathrm{r}}F(T^{+}\hat{T})+\frac{1}{\delta}E_{-,\mathrm{r}}K^{-1}F(T^{+}_{\lambda}\hat{T})\right)GE_{-,\mathrm{l}}-\left(\delta E_{-,\mathrm{r}}F(T^{+}\hat{T}_{\lambda})+E_{-,\mathrm{r}}K^{-1}F(T^{+}_{\lambda}\hat{T}_{\lambda})\right)GKE_{-,\mathrm{l}}\right),$
$\tilde{S}_{21}=\frac{1}{2}\left(\left(E_{+,\mathrm{r}}F(T^{+}\hat{T})-\frac{1}{\delta}E_{+,\mathrm{r}}K^{-1}F(T^{+}_{\lambda}\hat{T})\right)GE_{+,\mathrm{l}}+\left(\delta E_{+,\mathrm{r}}F(T^{+}\hat{T}_{\lambda})-E_{+,\mathrm{r}}K^{-1}F(T^{+}_{\lambda}\hat{T}_{\lambda})\right)GKE_{+,\mathrm{l}}\right),$
$\tilde{S}_{22}=\frac{1}{2}\left(\left(E_{+,\mathrm{r}}F(T^{+}\hat{T})-\frac{1}{\delta}E_{+,\mathrm{r}}K^{-1}F(T^{+}_{\lambda}\hat{T})\right)GE_{-,\mathrm{l}}-\left(\delta E_{+,\mathrm{r}}F(T^{+}\hat{T}_{\lambda})-E_{+,\mathrm{r}}K^{-1}F(T^{+}_{\lambda}\hat{T}_{\lambda})\right)GKE_{-,\mathrm{l}}\right).$
@@
where $\tilde{g}^{i}_{nm}:=2\mathrm{i}k^{(n)}g^{i}_{nm}$ and we use the notation $g^{i}:=\left(g^{i}_{nm}\right)_{m,n=-K}^{K}$. Here, $v_{i,m}^{\mathrm{in}}$ is the field that is incident to $D_{i}$, and the scattered field is defined piecewise by